OpenStudy (anonymous):

Please helpp. I am so confusedd. 1. You must match the number and the type of solution for the following two functions in standard form. f(x) = x^2 + 6x – 16 g(x) = x^2 +6x + 1 - Two real irrational solutions - Two real rational solutions

OpenStudy (anonymous):

Sorry, I think that will be in my class next year. :)

OpenStudy (anonymous):

I give medals. Please help!!

OpenStudy (anonymous):

@Isherron

OpenStudy (anonymous):

it is @Lsherron but lower case l

OpenStudy (anonymous):

ohh sorryt about that haha

OpenStudy (anonymous):

Do you get any of this?

OpenStudy (anonymous):

f(x) = x2 + 6x – 16 x1= 2 x2=-3 for g(x) = x2 +6x + 1 x1= −3+22√ x2= −3−22√

OpenStudy (anonymous):

please tell me that helped?

OpenStudy (anonymous):

look

OpenStudy (anonymous):

My teacher is super strict lol like im pretty sure he hates me, so is it at least close to right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you look at the pitcher?

OpenStudy (anonymous):

yess its cute lmaoo. do you think you can help me with another? this is a 3 part assignment.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did you getmy messeg?

OpenStudy (anonymous):

the next part is to explain how to convert f(x) to vertex form.

OpenStudy (anonymous):

i sent it